If theta is acute and (1-a square) sin theta = (1+a square) cos theta, then sin theta =

  • If theta is acute and (1-a square) sin theta = (1+a square) cos theta, then sin theta =

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

(1-a2) sinØ = (1+a 2) cos Ø
tanØ = (1+a2)/(1-a2)

So , sinØ = (1+a2)/ root{(1+a2)2+(1-a2)2}

Now simplify the above